# Accessible, At-Home Detection of Parkinson’s Disease via Multi-task Video Analysis

### v3

@@ Results
$F_{1}$ score AUROC AUPRC
@@
w/o MC-dropout $72.0\pm 0.9$ $69.0\pm 0.9$ $60.2\pm 1.3$ $73.9\pm 0.9$ $58.9\pm 1.0$
w/ MC-dropout $73.1\pm 0.7$ $70.1\pm 0.7$ $61.7\pm 0.9$ $74.9\pm 0.7$ $58.1\pm 0.9$
Smile
w/o MC-dropout $75.6\pm 0.2$ $72.2\pm 0.2$ $64.3\pm 0.3$ $83.2\pm 0.1$ $64.9\pm 0.2$
w/ MC-dropout $77.6\pm 0.2$ $74.4\pm 0.2$ $67.5\pm 0.3$ $83.6\pm 0.1$ $65.4\pm 0.1$
Speech
w/o MC-dropout $84.5\pm 0.3$ $82.5\pm 0.4$ $71.7\pm 0.4$ $\mathbf{89.4\pm 0.2}$ $\mathbf{81.9\pm 0.4}$
w/ MC-dropout $\mathbf{85.1\pm 0.2}$ $\mathbf{83.8\pm 0.5}$ $\mathbf{72.1\pm 0.6}$ $87.8\pm 0.1$ $80.7\pm 0.3$
Table 2: Performance of task-specific models with (w/) and without (w/o) MC-dropout (in percentage). Underlined metrics denote significantly better performance (within the same task), while bold metrics show overall best performance.
Among the three standardized tasks, the pangram utterance (speech) is the most accurate for classifying individuals with and without PD. Using only speech features, the shallow neural network achieved $84.5\pm 0.3\%$ accuracy and $89.4\pm 0.2\%$ AUROC. Finger-tapping had the least accuracy for PD detection. Models trained on right-hand tapping features performed better than those trained on left-hand features. However, the model trained on concatenated features from both hands significantly improved the $F_{1}$ score while being non-inferior in other metrics (see Technical Appendix B for details). For the remaining analysis, the finger-tapping model refers to the one trained on both-hand features. The model trained on smile features performed better than the finger-tapping task but worse than the speech task.
Effect of MC-dropout
MC-dropout significantly boosted the performance of the smile model in all metrics. For the speech model, while MC dropout improved accuracy and balanced accuracy, it decreased AUROC and AUPRC. The performance of the finger-tapping model was not notably affected. Due to the additional benefits of MC dropout in modeling prediction uncertainty, the multi-task models (UFNet and other baselines) were trained with MC-dropout unless specified otherwise. Detailed performance metrics of the task-specific neural networks along with the effect of MC dropout are reported in Table 2.
Effect of multi-task combinations
Combining multiple tasks using the proposed UFNet model enhanced performance (Table 3). For instance, the AUPRC scores of the multi-task models were significantly better than the corresponding single-task scores. Although the finger-tapping task alone was the weakest for detecting PD, its features may have complemented other task features, resulting in significant improvements in most metrics. Notably, combining all three tasks significantly improved all reported metrics, achieving an accuracy of $87.3\pm 0.3\%$, AUROC of $92.8\pm 0.2\%$, and an $F_{1}$ score of $81.0\pm 0.6\%$.
Task Combination Accuracy
@@
$\mathbf{F_{1}}$ score AUROC AUPRC
All three tasks $\mathbf{87.3\pm 0.4}$ $\mathbf{86.4\pm 0.4}$ $\mathbf{81.0\pm 0.6}$ $\mathbf{92.8\pm 0.2}$ $\mathbf{86.3\pm 0.5}$
Finger-tapping + Smile $78.0\pm 0.8$ $75.1\pm 1.0$ $65.6\pm 1.7$ $84.8\pm 0.5$ $74.5\pm 0.7$
Finger-tapping + Speech $84.1\pm 0.3$ $82.4\pm 0.3$ $77.3\pm 0.4$ $91.4\pm 0.2$ $84.5\pm 0.3$
Smile + Speech $85.2\pm 0.3$ $82.8\pm 0.5$ $75.0\pm 0.4$ $91.2\pm 0.1$ $82.2\pm 0.5$
Table 3: Performance of models (as percentages) trained on different combinations of the standardized tasks. The bold metrics denote best performance.
Model Accuracy

### v4

@@ Results
Task Accuracy $F_{1}$ score AUROC
Finger-tapping
w/o MC dropout $72.0\pm 0.9$ $60.2\pm 1.3$ $73.9\pm 0.9$
w/ MC dropout $73.1\pm 0.7$ $61.7\pm 0.9$ $74.9\pm 0.7$
Smile
w/o MC dropout $75.6\pm 0.2$ $64.3\pm 0.3$ $83.2\pm 0.1$
w/ MC dropout $77.6\pm 0.2$ $67.5\pm 0.3$ $83.6\pm 0.1$
Speech
w/o MC dropout $84.5\pm 0.3$ $71.7\pm 0.4$ $\mathbf{89.4\pm 0.2}$
w/ MC dropout $\mathbf{85.1\pm 0.2}$ $\mathbf{72.1\pm 0.6}$ $87.8\pm 0.1$
Table 3: Performance of task-specific models with (w/) and without (w/o) MC dropout (in percentage). Underlined metrics denote significantly better performance (within the same task), while bold metrics show overall best performance.
Among the three standardized tasks, speech is the most accurate for PD classification, and finger-tapping is the least accurate. The shallow neural network achieved $84.5\pm 0.3\%$ accuracy and $89.4\pm 0.2\%$ AUROC using only speech features. Models trained on right-hand tapping features performed better than those trained on left-hand features, indicating potential under-representation of left-handed individuals. However, the model trained on concatenated features from both hands significantly improved the $F_{1}$ score while being non-inferior in other metrics (see Appendix A). The finger-tapping model refers to the one trained on both-hand features for the remaining analysis.
MC dropout significantly boosted the performance of the smile model in all metrics. For the speech model, while MC dropout improved accuracy and balanced accuracy, it decreased AUROC and AUPRC. The performance of the finger-tapping model was not notably affected. Due to the additional benefits of MC dropout in modeling prediction uncertainty, the multi-task models (UFNet and other baselines) were trained with MC dropout unless specified otherwise. See Table 3 for details.
Effect of multi-task combinations
Combining multiple tasks using the proposed UFNet model enhanced performance (Table 4). For instance, the AUROC scores for all multi-task models were significantly better than the corresponding single-task scores. Although the finger-tapping task alone was the weakest for detecting PD, its features may have complemented other task features, resulting in significant improvements in most metrics. Combining all three tasks significantly improved all reported metrics, achieving an accuracy of $87.3\pm 0.3\%$, AUROC of $92.8\pm 0.2\%$, and an $F_{1}$ score of $81.0\pm 0.6\%$.
Task Combination Accuracy $\mathbf{F_{1}}$ score AUROC
All three tasks $\mathbf{87.3\pm 0.4}$ $\mathbf{81.0\pm 0.6}$ $\mathbf{92.8\pm 0.2}$
Tapping + Smile $78.0\pm 0.8$ $65.6\pm 1.7$ $84.8\pm 0.5$
Tapping + Speech $84.1\pm 0.3$ $77.3\pm 0.4$ $91.4\pm 0.2$
Smile + Speech $85.2\pm 0.3$ $75.0\pm 0.4$ $91.2\pm 0.1$
Table 4: Performance of models (as percentages) trained on different combinations of the standardized tasks. “Tapping” is used as the short form of the finger-tapping task. The bold metrics denote best performance.
Comparison against baselines
Model Accuracy
Balanced
Accuracy
AUROC $\mathbf{F_{1}}$ score Precision
Sensitivity
Specificity
Baseline models
Majority Voting 85.3 83.9 89.6 78.2 80.0 76.5 89.9
Neural Late Fusion $84.1\pm 0.4$ $81.3\pm 4.8$ $91.7\pm 2.2$ $73.2\pm 8.3$ $73.5\pm 7.5$ $76.3\pm 9.4$ $88.2\pm 2.7$
Early Fusion Baseline $83.6\pm 0.6$ $81.8\pm 0.7$ $91.0\pm 0.2$ $76.7\pm 0.7$ $75.4\pm 1.1$ $78.1\pm 0.9$ $86.5\pm 0.8$
Hybrid Fusion Baseline $84.1\pm 0.3$ $82.4\pm 0.4$ $91.4\pm 0.2$ $77.3\pm 0.4$ $76.2\pm 0.7$ $\mathbf{78.6\pm 0.6}$ $87.0\pm 0.6$
Our model with different attention modules
Dot product self-attention (Vaswani et al. 2017) $85.5\pm 0.4$ $84.3\pm 0.4$ $\mathbf{92.9\pm 0.2}$ $78.3\pm 0.6$ $80.7\pm 0.6$ $76.1\pm 1.1$ $90.4\pm 0.4$
LRFormer (Ye et al. 2023) $86.2\pm 0.5$ $85.1\pm 0.5$ $92.6\pm 0.3$ $79.5\pm 0.7$ $81.7\pm 0.9$ $77.6\pm 1.0$ $90.8\pm 0.6$
UFNet (ours) $\mathbf{87.3\pm 0.4}$ $\mathbf{86.4\pm 0.4}$ $92.8\pm 0.2$ $\mathbf{81.0\pm 0.6}$ $\mathbf{83.8\pm 0.5}$ $78.4\pm 1.0$ $\mathbf{92.0\pm 0.3}$
Early vs. hybrid fusion in UFNet
Early Fusion $86.7\pm 0.5$ $85.8\pm 0.5$ $92.7\pm 0.3$ $79.9\pm 0.8$ $83.3\pm 0.7$ $76.9\pm 1.4$ $91.9\pm 0.4$
Hybrid Fusion $\mathbf{87.3\pm 0.4}$ $\mathbf{86.4\pm 0.4}$ $92.8\pm 0.2$ $\mathbf{81.0\pm 0.6}$ $\mathbf{83.8\pm 0.5}$ $78.4\pm 1.0$ $\mathbf{92.0\pm 0.3}$
Table 5: Comparison of UFNet performance (as percentages) against multimodal baselines and other attention modules. The underlined metrics indicate significantly better performance than all four baselines. Bold metrics indicate the overall best performance across all choices. CI is not reported for majority voting since it does not involve any randomness.
The proposed uncertainty-calibrated fusion network (UFNet) outperformed the baseline methods in most metrics. More specifically, UFNet significantly improved PD classification accuracy, precision, and specificity compared to any of the four baselines. It also achieved higher balanced accuracy, AUROC, and $F_{1}$ score while maintaining statistically non-inferior sensitivity (Table 5).
Ablation results
To investigate the significance of our introduced uncertainty-calibrated self-attention module, we replace it with the standard dot product self-attention (Vaswani et al. 2017) and LRFormer (Ye et al. 2023) – a recently proposed self-attention module that prevents transformer overconfidence via Lipschitz Regularization. In addition, UFNet uses task-specific features and prediction scores, making it a hybrid fusion approach. We implement an early fusion approach to analyze the effect of removing task-specific predictions from UFNet.
The calibrated self-attention module outperformed the other two attention modules in all metrics except AUROC. While the dot product self-attention yielded 0.1% higher (non-significant) AUROC, UFNet was significantly better in most other metrics. LRFormer performed closely in most metrics. Finally, including task-specific predictions as additional input (hybrid fusion) enhanced performance over the early fusion approach. Since our task-specific models were trained with more subjects, they likely better understand task-specific predictions, and incorporating them improved final performance.
Withholding uncertain predictions
The trivial usage of MC dropout improved prediction accuracy. With dropped uncertain predictions, the best UFNet (hybrid fusion) model achieved $88.0\pm 0.3\%$ accuracy, $87.1\pm 0.3\%$ balanced accuracy, $93.0\pm 0.2\%$ AUROC, and $81.8\pm 0.5\%$ $F_{1}$ score – a clear improvement over our base model. It also achieved better precision, sensitivity, and specificity. However, the model could now predict $97.8\pm 0.3\%$ of the data where it is confident enough. The model’s expected calibration error (ECE) was $5.4\pm 0.5\%$, and the Brier score was $0.097\pm 0.002$, indicating that the model-predicted probabilities aligned well with actual disease probabilities.
Withholding Approach Accuracy Precision Recall F1 Score AUROC Coverage
MC Dropout + 95% CI $88.0\pm 0.3$ $84.6\pm 0.5$ $79.3\pm 0.9$ $81.8\pm 0.6$ $93.0\pm 0.2$ $\mathbf{97.8\pm 0.3}$
Conformal Prediction $93.9\pm 0.7$ $87.0\pm 1.4$ $89.0\pm 2.5$ $87.9\pm 1.8$ $93.9\pm 0.7$ $55.1\pm 2.5$
Label Smoothing + Conformal Prediction $92.3\pm 0.7$ $84.9\pm 1.2$ $90.0\pm 1.5$ $87.3\pm 1.0$ $94.4\pm 0.6$ $56.9\pm 2.4$
Platt Scaling + Conformal Prediction $91.9\pm 0.8$ $85.3\pm 0.9$ $\mathbf{97.6\pm 0.7}$ $\mathbf{91.0\pm 0.4}$ $92.8\pm 2.2$ $48.7\pm 3.3$
Label Smoothing +
Platt Scaling + Conformal Prediction
$\mathbf{94.2\pm 0.5}$ $\mathbf{87.4\pm 1.3}$ $92.9\pm 1.3$ $89.9\pm 0.7$ $\mathbf{95.5\pm 0.7}$ $52.0\pm 2.6$
Table 6: Different methods for withholding predictions.
Applying conformal prediction further reduced misprediction at the cost of a more frequent prediction withholding. Simultaneously applying label smoothing, Platt scaling, and conformal prediction resulted in a notable improvement in all metrics, resulting in $94.2\pm 0.0.5\%$ accuracy, $89.9\pm 0.7\%$ $F_{1}$ score, and $95.5\pm 0.7\%$ AUROC. But this model could predict with enough certainty in only $52.0\pm 2.6\%$ cases, throwing away approximately half of the test data. Please see Table 6 for details.
Performance across demographic subgroups
No significant bias in model performance was observed (see Figure 2) based on sex or race in the test set (162 individuals). The average error (i.e., misclassification) rate across female participants ($n=85$) was $14.1\pm 7.4\%$, compared to $6.5\pm 5.5\%$ for male participants ($n=77$). This difference is notable but not statistically significant (p-value = 0.11 using two-sample $Z$-test for proportions). The error rate was $7.63\pm 4.79\%$ for white participants ($n=118$) and $5.56\pm 11.39\%$ for non-white participants ($n=18$). This difference was also non-significant based on Fisher’s exact test (Fisher’s odd ratio = 0.71, p-value = 1.0). However, the error rate varied notably based on age subgroups. The model performed well for individuals aged 50 to 80. At the same time, the error rate was higher for those aged 30 to 50 or over 80, likely due to the under-representation of these age groups, as the dataset primarily consists of individuals who are 50-80 years old (77.1% of the entire dataset).
Figure 2: Misclassification rate of the best UFNet model across demographic subgroups. The analysis is done on the test set participants. Error bars demonstrate 95% CIs.
Validation with external datasets
Since we introduced the first multi-task dataset for PD detection in this study, we could not compare our model to similar external datasets. However, we experimented with our model on a recently released public dataset named YouTubePD (Zhou et al. 2024). This dataset consists of 283 short YouTube videos collected from 200+ individuals (16 confirmed PD patients). Even without adapting the feature extraction pipeline for this new task, our model (UFNet with Smile + Speech modalities) performed comparably to state-of-the-art – 18% improvement in accuracy, 3% reduction in AUROC and $F_{1}$ score (see Table 7). This result is promising as our video features are designed for smile and pangram utterance tasks, vastly different from the videos in the YouTubePD dataset. Adapting the features for this new task should further improve performance.
Model Accuracy $\mathbf{F_{1}}$ score AUROC
YouTubePD (Zhou et al. 2024) $70.61\%$ $\mathbf{61.0\%}$ $\mathbf{87.0\%}$
UFNet (Smile + Speech) $\mathbf{88.76\%}$ $58.33\%$ $83.83\%$
Table 7: Summary of performance on the YouTubePD dataset. The bold metrics denote best performance.
Discussions and Limitations
Video analysis provides an accessible, cost-effective, and convenient means of screening for PD, particularly benefiting individuals in remote areas or low-income countries with limited access to neurological care. We carefully selected three proposed tasks for remote completion, considering feasibility and safety. While gait analysis is standard for evaluating PD (Li and Li 2022; Liu et al. 2022), it presents logistical challenges and potential risks of falling for PD patients. Sustained phonation (holding a vowel sound for as long as possible) is another speech assessment option (Tsanas et al. 2009; Vaiciukynas et al. 2017), but its analysis becomes complicated due to inconsistencies across various recording devices. All selected tasks demonstrated reasonable predictive performance in differentiating individuals with and without PD, validating our task selection.
Despite our model’s promise of accurate PD detection, harm from misprediction remains a concern. Falsely classifying someone as a PD patient may cause undue stress and economic burden, while missing individuals with PD provide a false sense of security and create delays in getting clinical care. With 78.4% recall, our model may miss some individuals with PD, so we recommend clinical follow-up if symptoms persist or worsen, even if our model does not detect PD. We also advise individuals not to consider our model’s positive prediction a definitive diagnosis. Finally, we reduce the harm of mispredictions by withholding predictions when the model is uncertain.
Our model performs consistently across sex and ethnic subgroups, but accuracy drops for younger (30-49 years) and older (over 80 years) age groups. We find that participants from these groups are under-represented, as over 75% of our dataset participants fall within the 50-80 age range. Until a more balanced dataset is available, we recommend applying the tool only to individuals between 50 and 80. Also, due to requiring an English pangram utterance, our best model cannot be used for non-English speakers. Future work should prioritize recruiting more participants and investigating speech tasks in other languages for the worldwide deployment of our methods.
The decision threshold used in the study could be customized based on individual preferences, as the it directly affects the model’s sensitivity and specificity. We used a common 0.5 threshold, but individual preferences for risk-benefit trade-offs might necessitate adjustments. For instance, some users might seek clinical evaluation even at lower probabilities (i.e., prefer high sensitivity), while others might wait for a higher likelihood before incurring healthcare costs (i.e., prefer high specificity).
Furthermore, as videos are primarily gathered unsupervised, issues such as noncompliance with task instructions and various forms of noise are common. For example, during the finger-tapping task, many participants performed fewer than the required ten taps, often with their task-performing hand obscured from view. Background noise may distort speech features, and the presence of multiple individuals in the frame could compromise smile feature extraction. Integrating post hoc quality assessment algorithms into the data collection framework could enhance data quality by identifying quality issues and prompting users to re-record tasks if needed.
A comprehensive, large-scale dataset for Parkinson’s disease is currently unavailable due to the lack of resources and clinical collaboration. This paper addresses this gap by introducing the largest multi-task video dataset in the literature and publicly releasing the anonymized features. Unfortunately, we could not gather multiple datasets of videos or extract similar features to compare model performance across different datasets since patient videos (protected health information) are not publicly accessible. Future research on privacy-preserving ways of sharing video datasets can help bridge this gap, facilitating broader comparisons and furthering research in health AI.
Conclusion
Acknowledgments
The project was supported by the National Science Foundation Award IIS-1750380, National Institute of Neurological Disorders and Stroke of the National Institutes of Health under award number P50NS108676, and Gordon and Betty Moore Foundation. The first author (MSI) is supported by a Google PhD fellowship.
Appendix A. Supplementary Results
Co-variate analysis of demographic variables
To identify whether there is any sampling bias in our dataset splits, we investigate the distribution of the various demographic attributes across the splits. We found that individuals from different demographic groups are identically distributed across training, validation, and test splits (see Table 8).
Number of Individuals from the Minority Class (%)
Demographic Attribute Classes Minority Class Train (n = 516) Validation (n = 167) Test (n = 162)
Health Condition with PD, without PD with PD 168 (32.6%) 55 (32.9%) 49 (30.3%)
Sex* Male, Female Male 243 (47.1%) 80 (47.1%) 77 (47.5%)
Age Below 50, Above 50 Below 50 48 (9.3%) 14 (8.4%) 17 (10.5%)
Ethnicity white, Non-white Non-white 44 (8.5%) 13 (7.8%) 18 (11.1%)
Recording Environment Home, Non-home Non-home 193 (37.4%) 72 (43.1%) 54 (33.3%)
Table 8: Co-variate analysis. Non-binary sex group is not used for simplicity of presentation.
Performance of hand-separated finger-tapping models
Individuals with PD may experience asymmetry in their motor performance at the onset of the disease. Therefore, to model the finger-tapping task, that assess motor performance, it may be beneficial to utilize both hands’ features simultaneously. To investigate this, we trained three task-specific models separately (without employing Monte Carlo dropout). We report their performance in Table 9, and observe that utilizing features from both hands significantly improves the $F_{1}$ score of the model, while maintaining non-inferiority across other metrics.
@@ Appendix A. Supplementary Results
$\mathbf{F_{1}}$ score AUROC AUPRC
@@
Both hands
72.0
[71.1, 72.9]
69.0
[68.1, 70.0]
60.2
[58.9, 61.5]
73.9
[73.1, 74.8]
58.9
[57.9, 59.8]
Left hand
64.3
[62.9, 65.6]
62.0
[60.7, 63.4]
52.6
[50.8, 54.5]
66.9
[65.3, 68.6]
51.8
[50.2, 53.4]
Right hand
73.0
[72.3, 73.7]
70.0
[68.9, 71.1]
47.3
[44.4, 50.2]
73.7
[72.5, 75.0]
58.5
[57.0, 60.0]
Table 9: Performance of the task-specific finger-tapping models. Underlined metrics show significantly better performance.
Performance of pre-trained video models
We experimented with features obtained from four pre-trained video models (VideoMAE (Tong et al. 2022), Uniformer (Li et al. 2023), ViViT (Arnab et al. 2021), and TimeSformer (Bertasius, Wang, and Torresani 2021)) for classifying PD based on the finger-tapping and smile tasks, and compared them against the custom features used in this study. The performance on the finger-tapping task is detailed in Table 10, and the smile task in Table 11.
Feature Extraction Method Accuracy
@@
F1 Score AUROC AUPRC
Custom features $\mathbf{73.1\pm 0.7}$ $\mathbf{70.1\pm 0.7}$ $\mathbf{61.7\pm 0.9}$ $74.9\pm 0.7$ $58.1\pm 0.9$
VideoMAE $70.2\pm 0.5$ $66.8\pm 0.5$ $57.0\pm 0.9$ $75.1\pm 0.3$ $60.6\pm 0.5$
Uniformer $71.9\pm 0.4$ $68.3\pm 0.6$ $54.5\pm 1.6$ $74.3\pm 0.6$ $55.6\pm 0.8$
ViViT $71.6\pm 0.5$ $68.0\pm 0.6$ $50.8\pm 1.9$ $73.7\pm 0.4$ $57.5\pm 0.5$
TimeSformer $72.6\pm 0.3$ $68.9\pm 0.4$ $55.9\pm 1.6$ $\mathbf{77.2\pm 0.3}$ $\mathbf{61.3\pm 0.5}$
Table 10: Finger-tapping task performance
Feature Extraction Method Accuracy